solve by the elimination method p=40+q and 7p=-9q+360
Answer by TimothyLamb(4379)

---
p = 40 + q
7p = -9q + 360
---
put the system of linear equations into standard form:
---
x = p
y = q
---
x = 40 + y
7x = -9y + 360
---
x - y = 40
7x + 9y = 360
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution by matrix method:
x = p = 45
y = q = 5
---
now by elimination:
---
x - y = 40
7x + 9y = 360
---
p - q = 40
7p + 9q = 360
---
( p - q = 40 ) * -7
7p + 9q = 360
---
-7p + 7q = -280
7p + 9q = 360
---
add the system:
---
16q = 80
q = 5
---
p = 40 + q
p = 40 + 5
p = 45
---
solution by elimination method:
x = p = 45
y = q = 5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php